OpenStudy (anonymous):

Anyone can help me to check my answer(in my first reply)? Thanks For any event A, A is a subset of the sample space S. show that E(IA) = P(IA = 1) = P(A). IA is Indicator random variable

OpenStudy (anonymous):

If \[A\] is event occurs, \[I_{A} = 1\] So \[I_{A} = 0\] is event Not occurs, then \[E(Y) = \sum_{y} y P(Y=y) | y=0,1 \] \[E(Y) = 0P(Y=0) + 1P(Y=1) = P(Y=1)\] As we know from Bernoulli distribution \[P(I_{A}=1) = P(A) \\and\\ P(I_{A}=0) = 1 - P(A)\] Then \[E(I_{A}) = P(I_{A}=1) = P(A)\]

OpenStudy (anonymous):

Hi @dumbcow , Sorry to bother you again, could you please help me check my answer again for this new question?

OpenStudy (dumbcow):

everything looks correct to me...you may want a 2nd opinion just to confirm it

OpenStudy (anonymous):

Thanks @dumbcow, you are really helpful. I will leave this open to see if others would like to comment.

OpenStudy (anonymous):

looks good to me!

OpenStudy (anonymous):

Thanks @pgpilot326 , I can close this question now :)